# Coherent Source Subsampling: A Data-Driven Strategy for Restoring Causal–Acausal Symmetry in Ambient Seismic Wavefield Correlations

### v1

@@ 3 Results
Figure 2: Cross-correlation gathers for the eastern United States. (a) Before subsampling: The linearly averaged cross-correlations, $\widehat{\mathbb{E}}\left[\prescript{b}{}{\mathbf{w}}_{ij}\right]$, show large amplitudes close to zero lag and pronounced causal-acausal asymmetry, indicating a directionally biased ambient noise field. (b) After subsampling: The conditional averages, $\widehat{\mathbb{E}}\left[\prescript{b}{}{\mathbf{w}}_{ij}\mid\lambda=1\right]$, obtained after subsampling, display enhanced waveform coherence and substantially improved causal–acausal symmetry. The percentages of time-window allocation (in red) indicate the fraction of windows selected for the source state $\lambda\text{ = 1}$. The correlation values (in blue) measure the degree of symmetry between the acausal and causal components for each station pair. The black lines mark the time lags corresponding to velocities of $2$ and 55\,km/s.
[…]
The improvement in cross-correlation symmetry translates directly into enhanced Rayleigh-wave dispersion measurements. Figures 3c–e for (GOGA, NCB) and 3k–m for (BLA, PKME) display the Rayleigh-wave dispersion images obtained without subsampling, which show asymmetry between the causal and acausal branches. After applying CSS averaging for $\lambda=1$, the dispersion images (Figures 3f–h and 3n–p) exhibit markedly enhanced coherence and symmetry. This improved balance between the causal and acausal branches enables a much more reliable picking of group velocities. Figures 4a–i show group-velocity dispersion curves, extracted by automated picking, for nine station pairs with short, intermediate, and long interstation distances. For each pair, the dispersion curves obtained by simple linear averaging display substantial scatter and a clear mismatch between the causal and acausal branches. After applying conditional averaging, the dispersion estimates become much more stable: the causal and acausal branches closely coincide, producing nearly symmetric behavior in the 5505–50\,s period band, where Rayleigh-wave energy is strongest. The pairs with the longest and intermediate distances, including (GOGA, NCB) (14201420\,km), (BLA, PKME) (12901290\,km), (CBN, PKME) (10321032\,km), and (ACSO, CNNC) (713713\,km), show the most pronounced improvements, with the subsampled curves collapsing onto a well-defined trend. Other station pairs, such as (AAM, GOGA) (989989\,km), (BINY, GLMI) (755755\,km), (AAM, BINY) (632632\,km), (ACSO, BINY) (625625\,km), and (BLA, GOGA) (505505\,km), also exhibit notably closer agreement between the causal and acausal branches than in the linear averages. Finally, we compare the CSS-derived dispersion curves with the earthquake-derived dispersion curves for (GOGA, NCB) and (ACSO, CNNC) pairs, using earthquakes of about $M_{w}\,4.5$ located near the GOGA and ACSO stations, respectively. For both pairs, the subsampled symmetric dispersion curves agree well with the earthquake-based dispersion curves.
It is important to note that while CSS enhances time-domain symmetry, it does not guarantee identical causal and acausal dispersion curves across all frequency bands. This is an expected consequence: if the stationary zone lacks sources at specific frequencies, or if certain frequency bands are naturally weak or absent in the ambient field, CSS can enhance the coherence of existing energy but cannot artificially generate it. Thus, the method optimizes the extraction of information from the available ambient field but is fundamentally constrained by the spectral content of that field.
Figure 3: Temporal and dispersion analysis for the station pairs (GOGA, NCB) and (BLA, PKME). (a–b, i–j) Acausal and causal time-window selections obtained from CSS for source states $\lambda\text{ = \{1, 2\} }$. (c–e, k–m) Before subsampling: standard linear averaging produces dispersion images and cross-correlations with low coherence and pronounced causal–acausal asymmetry. (f-h, n-p) After subsampling: source state $\lambda\text{ = 1}$ results in dispersion images and cross-correlations with substantially improved causal–acausal symmetry.
Figure 4: Group‐velocity dispersion curves for nine station pairs in the eastern United States. Solid lines represent the causal and acausal dispersion curves obtained for source state $\lambda\text{ = 1}$, while dotted lines indicate the corresponding dispersion curves based on linear averaging. Earthquake‐based dispersion curves are shown for comparison for station pairs (GOGA, NCB) and (ACSO, CNNC). For all station pairs, the dispersion curves retrieved using source state $\lambda\text{=1}$ show markedly reduced scatter between 1010\, and 5050\,s, whereas the linearly averaged curves display substantial variability and systematically underestimate velocities. This closer agreement demonstrates the effectiveness of CSS for recovering symmetric dispersion information.
@@ 4 Discussion
The results show that CSS effectively isolates the relatively small fraction of ambient-noise windows that exert the strongest control on empirical Green’s function retrieval, yielding substantial improvements in causal–acausal symmetry and dispersion stability. By extracting quasi-stable noise regimes that recur over long periods, CSS provides a physically interpretable connection between ambient-noise variability and the theoretical conditions required for reciprocal Green’s function estimation.
Because it reduces the number of windows needed to achieve coherent stacking, CSS is particularly well suited to large-scale experiments, including dense nodal arrays and distributed acoustic sensing (DAS). In such data-heavy environments, CSS can retrieve coherent surface-wave energy even in the presence of intense anthropogenic noise and pronounced spatial heterogeneity, enhancing interferometric quality while greatly decreasing data volume.
More broadly, the conditional-averaging framework at the core of CSS may benefit applications where coherent wavefield reconstruction is constrained by uncertain or varying source distributions. A current practical limitation is that our implementation trains a separate VAE for each station pair; while this is feasible for modest-sized networks, scaling to hundreds of pairs would require substantial computational effort or new joint, multi-pair training strategies. Developing such scalable approaches is a crucial step toward deploying CSS at network-wide or continental scales. In future work, CSS could also be adapted to independently subsample narrow frequency bands in order to detect frequency-specific weak sources within the stationary zone.
@@ 5 Conclusion
We introduced Coherent Source Subsampling (CSS), a data-driven approach that reformulates ambient noise averaging as a conditional process guided by source-state coherence. Instead of relying on indiscriminate stacking, CSS computes conditional averages over subsets representing quasi-stationary segments of the ambient field. By isolating these coherent intervals, CSS suppresses the influence of transient and directionally biased noise and substantially enhances the causal–acausal symmetry of the retrieved cross-correlations.
Application to the eastern United States dataset demonstrates that coherent Green-function estimates can be obtained from only a small fraction of all available time windows. Conditional averages produced through CSS yield clearer Rayleigh-wave arrivals, improved causal–acausal agreement, and more stable group-velocity dispersion curves across a range of interstation distances. These improvements reduce the bias and variability inherent in linear stacking and provide dispersion measurements better suited for robust ambient-noise tomography. In general, CSS offers a physically interpretable and empirically validated pathway for extracting reliable structural information from nonstationary and anisotropic noise fields, extending the applicability of ambient noise imaging in challenging environments.
@@
Waveform data used in this study can be obtained from the IRIS EarthScope Data Management Center (https://ds.iris.edu/ds/nodes/dmc/) through the FDSN dataselect service (http://service.iris.edu/fdsnws/dataselect/1/). The dataset includes recordings from the United States National Seismic Network (network code US) USArray. All codes developed as part of this work will be made publicly available after the manuscript undergoes peer review.
@@ Acknowledgments
This work was supported by the Shell project under project titled Neural Network Based Denoising Of Passive Seismic Data. We would like to thank Dr. Anu Chandran (Shell), Dr. Rene-Edouard Plessix (Shell), Dr. Arjun Datta (IISER Pune), Prof. Shyam Rai (IISER Pune), Prof. Aurélien Mordret (GEUS), Prof. Gerard Schuster (University of Utah) and Dr. Kees Weemstra (KNMI) for their valuable feedback. We thank Tiente Rengneichuong Koireng (IISc) and Pushkar Bharadwaj (IISc) for testing the method on other datasets. The authors used AI-assisted tools (OpenAI ChatGPT and Writefull) for language refinement. All edits were reviewed, and the authors are solely responsible for the final text.

### v2

@@ 3 Results
Figure 2: Cross-correlation gathers from the Central California array. (a,c) Linear averaging without subsampling, $\widehat{\mathbb{E}}\left[\prescript{b}{}{\mathbf{w}}_{ij}\right]$, for the 3388\,s and 885050\,s bands, respectively. The gathers exhibit clear causal–acausal asymmetry, with higher signal-to-noise ratios on the acausal side due to dominant illumination from Pacific Ocean microseisms. (b,d) The conditional averaging, $\widehat{\mathbb{E}}\left[\prescript{b}{}{\mathbf{w}}_{ij}\mid\lambda=1\right]$ using CSS for the same bands, shows improved symmetry between causal and acausal branches. Red percentages denote the fraction of time windows assigned to the source state $\lambda=1$ by CSS, and blue values show the correlation between causal and acausal branches for each station pair. The black lines mark time lags corresponding to apparent velocities of $1.25$ and 55\,km/s.
For the Central California array, the correlations after averaging without subsampling, $\widehat{\mathbb{E}}\left[\prescript{b}{}{\mathbf{w}}_{ij}\right]$, shown in Figures 2a and 2c, reveal a strongly directional ambient noise wavefield and display clear causal–acausal asymmetry. For both 3388\,s and 885050\,s period bands, the dominant energy is present in the acausal branch, consistent with seismic energy primarily originating from the Pacific Ocean. Because the acausal branch exhibits a signal-to-noise ratio (SNR) much higher than that of the causal branch, it is typically used for tomographic imaging. However, it should be noted that this measurement may be biased relative to the true interstation response, as indicated by the observed asymmetry. In contrast, the conditionally averaged correlations (subsampled averages), $\widehat{\mathbb{E}}\left[\prescript{b}{}{\mathbf{w}}_{ij}\mid\lambda=1\right]$, shown in Figures 2b and2d exhibit a markedly stronger symmetry: the Rayleigh waves on the causal and acausal branches are nearly identical. This enhanced symmetry is quantified in Figure 2. For each pair of stations, Figure 2 also reports the fraction $\frac{\left|\prescript{b}{\lambda}{\mathcal{C}}_{ij}\right|}{M_{ij}}$ of the retained time windows. The increased causal–acausal similarity indicates that CSS preferentially selects windows associated with the stationary source illumination state. We now obtained correlations that more accurately represent the true interstation response.
In Figure 3, we show that the enhanced symmetry improves the dispersion measurements. Here, the matrices derived without subsampling display pronounced artifacts: the group velocities vary irregularly between adjacent station pairs and across periods, particularly for the low-SNR branch (upper triangles of the matrices in Figures 3a, c and e). Although the high-SNR group-velocity picks (lower triangles of Figures 3a, c and e) do not exhibit significant artifacts, they may still be influenced by bias arising from non-uniform source illumination. Our subsequent tomography findings will further support the presence of this bias. In contrast to linear averaging, we observe that the CSS group-velocity picks are more similar across the branches — note that the artifacts in the low-SNR branch are significantly reduced. In addition, the velocity picks in the high-SNR branch change slightly before and after subsampling, which reflects the reduction in bias.
Consequently, we construct three sets of group‐velocity tomograms (see Figure 4) using: 1. picks from the high-SNR branch without subsampling — hereafter, referred to as LS-high tomograms; 2. picks obtained without subsampling but after averaging the high- and low-SNR branches — hereafter, referred to as LS tomograms; 3. picks derived from the correlations after CSS. For CSS tomography, since the group-velocity picks are generally consistent across the different branches, we averaged the values when they matched within a 15% range and discarded the picks otherwise. We used the surface‐wave tomography module of SeisLib (Magrini et al., 2022), which performs a regularized least-squares tomography of interstation group‐velocity measurements. The regularization parameters were selected to yield comparable apparent smoothness and resolution in all cases so that the differences discussed below arise from the group-velocity picks themselves rather than from inversion settings.
  • Reduction of structural bias relative to fault geometry: In the LS and LS-high models, velocity anomalies are elongated along the station array and approximately parallel to the dominant raypaths rather than to the mapped faults. Such along-array smearing reflects the bias due to non-uniform source distribution. In contrast, the CSS tomograms show velocity contrasts that follow the tectonic structure: distinct low-velocity zones adjacent to the RF and the SAF fault system, and a relatively higher-velocity block confined between them. This behavior is consistent with controlled-source and local-earthquake imaging studies near Parkfield, which document reduced velocities within the fault damage zone and comparatively higher velocities in adjacent crustal blocks (Hole et al., 2006; Yang et al., 2025) (Supplementary Figure S77). The improved alignment of anomalies with mapped structures indicates that CSS reduced source illumination bias and better recovers geologically consistent heterogeneity.
  • Improved consistency between periods and physical plausibility: Surface-wave group velocities should evolve smoothly with period because sensitivity kernels broaden gradually with depth. The CSS models exhibit this expected behavior: the inter-fault block persists from  88\,s to 11~11\,s (roughly 10101313\,km depth) and weakens progressively with increasing period. By contrast, LS models display abrupt appearance and disappearance of anomalies and inconsistent velocity changes between neighboring periods. Such an inconsistency is unlikely to reflect real structure and instead suggests contamination from period-dependent source distribution, a known limitation of ambient-noise correlation under non-uniform illumination. Note that applying smoothing to reduce these inconsistencies will lead to a loss of resolution.
  • Shallow fault‐zone effects: In shorter periods, CSS tomograms consistently resolve lower group velocities along the RF and SAF fault traces, while the intervening block remains faster. This pattern matches the expected seismic signature of a fault damage zone: fractured and fluid-altered rocks reduce seismic velocities within the fault core and surrounding damage region, while more intact crystalline rocks between faults maintain higher velocities. Similar low-velocity fault-zone structures have been reported from borehole and seismic studies at SAFOD and related Parkfield experiments (Hole et al., 2006). The LS tomograms either smear these regions along the array or fail to isolate them.
These observations demonstrate that CSS improves ambient noise tomography by recovering structurally coherent and geologically interpretable features. We also processed another ambient-noise dataset (Liang and Langston, 2008) from the eastern United States, a stable cratonic region, to provide additional evidence; these results are presented in the Supplementary Text S44.
Figure 3: Group‐velocity matrices obtained from ambient noise cross-correlations. For each subplot, the columns correspond to virtual source stations and rows to receiver stations (CC0101–CC2020). The color scale indicates Rayleigh-wave group velocity (km/s). Top row (a,c,e) shows picks without subsampling for periods of 88\,s, 99\,s, and 1010\,s, respectively, while the bottom row (b,d,f) shows the corresponding picks after Coherent Source Subsampling (CSS). White regions represent the lack of picks e.g., due to a smaller offset relative to Rayleigh-wave wavelength. The gray diagonal annotation reports the average similarity between the causal and acausal branch picks. Compared to linear averaging, CSS yields significantly higher reciprocity.
Figure 4: Group‐velocity tomograms (a–l) and ray coverage (m) for the Central California array, where the stations (red inverted triangles) are roughly along the transect marked using the dashed magenta line. To obtain the tomograms for periods of 881111\,s (approximate peak sensitivity at depths of  10101313\,km) the group-velocity picks are derived from: the high-SNR branch only (left column), both high- and low-SNR branches (middle column), and the proposed CSS selection (right column). Major tectonic elements, including the Southern Coast Ranges (SCR), Rinconada Fault (RF), San Andreas Fault (SAF), and Great Valley Basin (GVB), are indicated. The CSS tomograms mirror the underlying geology and reveal a high-velocity block located between the RF and SAF, a feature that is consistently supported by previous studies based on either refraction data or local-earthquake traveltimes.
@@ 4 Discussion
As CSS can effectively extract interstation response, it is particularly well suited for experiments using dense nodal arrays and distributed acoustic sensing (DAS). In these data-intensive settings, the goal of interferometry is to reduce the data volume while preserving useful subsurface information. CSS can potentially be used to mitigate biases due to strong directional anthropogenic noise sources in these settings, improving interferometric extraction of interstation response.
The current practical limitation is that our implementation trains a separate VAE for each station pair; while this is manageable for arrays of moderate size, extending to hundreds of pairs would demand substantial computational resources. Development of new, joint multi-pair training schemes is essential as such scalable techniques are a key step toward applying CSS to dense seismic surveys.
@@ 5 Conclusion
We introduced Coherent Source Subsampling (CSS), a data-driven approach that reformulates ambient noise interferometry as a conditional averaging process guided by a neural-network based stationary-zone source selection. CSS suppresses the influence of transient and directionally biased noise and substantially enhances the causal–acausal symmetry of the retrieved cross-correlations. Application to the central California dataset demonstrates that subsampling yields more reliable Rayleigh-wave arrivals, improved causal–acausal agreement, and more stable group-velocity dispersion curves across a range of interstation distances. These improvements reduce the bias and variability inherent to traditional ambient-noise tomography done without subsampling. We demonstrated that the CSS tomograms successfully image established geological structures, including a high-velocity block located between the Rinconada and San Andreas faults, as well as low-velocity zones adjacent to these faults. In summary, CSS extends the applicability of ambient noise interferometry to challenging environments with complex source distributions and shorter survey durations.
@@
Waveform data used in this study can be obtained from the IRIS EarthScope Data Management Center (https://ds.iris.edu/ds/nodes/dmc/) through the FDSN dataselect service (http://service.iris.edu/fdsnws/dataselect/1/). The dataset includes recordings from the CCSE array (network code TO) (CCSE, 2015) and United States National Seismic Network (network code US) (Albuquerque Seismological Laboratory (ASL)/USGS, 1990). The codes developed for this study are included as a ZIP archive in the filetype “Dataset” for peer review and will be made publicly available upon publication.
@@ Acknowledgments
This work was supported by the Shell project under the project titled Neural Network Based Denoising Of Passive Seismic Data. We would like to thank Dr. Anu Chandran (Shell), Dr. Rene-Edouard Plessix (Shell), Dr. Arjun Datta (IISER Pune), Prof. Shyam Rai (IISER Pune), Prof. Aurélien Mordret (GEUS), Prof. Gerard Schuster (University of Utah), and Dr. Kees Weemstra (KNMI) for their valuable feedback. We also thank Prof. Brandon Schmandt (Rice University) for discussions on the CCSE dataset. We thank Tiente Rengneichuong Koireng (IISc) for testing the method on other datasets. The authors used AI-assisted tools (OpenAI ChatGPT and Writefull) for language refinement. All edits were reviewed, and the authors are solely responsible for the final text.